### v1

Abstract
Capital growth, at large scales only, arrives with no help from net saving, and consequently with no help from consumption constraint. Net saving, at large scales, is sacrifice of consumption with nothing in return.
keywords:
[…]
1 Introduction
Our "Sources of capital growth" [^1] Getty and Tkachenko (2024) proposed the term "thrift theory" to mean the commonly accepted doctrine that net saving, within the technological growth rate ("warranted rate"), gives capital growth. "Free growth theory" was defined as an opposite hypothesis that capital growth is explained wholly by a gain in productivity of capital and labor already present, so that net saving is superflous. Data for net saving and market-value capital taken from natural accounts of 92 nations were argued to support free growth theory. Here we pursue a parallel argument and test from data for consumption.
2 Consumption and net saving
Thrift theory begins with the idea that net saving $S_{net}$, within the warranted rate, gives capital growth $\Delta K$. That is,
$\Delta K=S_{net}\ ,\quad\text{in thrift theory.}$. (1)
where "within the warranted rate" is implicit. We divide this by capital (normalize) to avoid nonsense correlation when finding averages over time. Define $g(K)=\frac{\Delta K}{K}$ and $s^{*}=\frac{S_{net}}{K}$ to infer, from Eq. (1),
$g(K)=s^{*}\quad\text{in thrift theory.}$. (2)
In thrift theory and common teaching, where innovation provides a required condition for capital growth, but only insofar as net income is saved in investment, and where consumption and net saving sum to net income; it follows that any change in net saving implies an equal and opposite change in consumption $C$222i.e., a suspension of consumption provides the only source of a gain in net saving, arguendo, while a gain in consumption becomes the only disposition of net income not applied to net saving. That is,
$\Delta S_{net}=-\Delta C\ ,\quad\text{in thrift theory.}\quad\text{Define }c^{*}=\frac{C}{K}\text{ to proceed with}$ (3)
$\Delta g(K)=\Delta s^{*}=-\Delta c^{*}\ ,\quad\text{from which}\quad\frac{-\Delta c^{*}}{\Delta g(K)}=1\ ,\quad\text{in thrift theory.}$. (4)
For convenience, define
[…]
$\theta$ in general is what we call the "thrift index," which measures the success of the predictions of thrift theory in Eqs. (4) and (5) against a standard of unity (the number one), and $\theta_{c}$ does so indirectly from data for consumption. [^4] An argument from data for net saving and consumption is offered in Getty and Tkachenko (2024). Displays below give observations of $\theta_{c}$ from data for $C$ and market-value capital $K$ in national accounts in 90 countries over the period 1980-2021.
Fig. 1: Plots of $\theta_{c}$ averaged over all countries, GDP-weighted. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\ \theta_{c}=1$.
Table 1: Regression of $-\Delta c^{*}$ on $\Delta g(K)$, all countries and years, GDP-weighted. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\text{Regression}=1$.
Regression -0.129***
Std. error 0.021
Observations 772
$R^{2}$ 0.404
Within $R^{2}$ 0.18
Year fixed effects Yes
Country fixed effects Yes
***$p<0.01$, **$p<0.05$, *$p<0.1$
Table 2: $\theta_{c}$ in Countries and Periods Shown. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\theta_{c}=1.$
Country $\theta_{c}$ Periods
Armenia 0.29 97-19(11)
Aruba 1.85 00-00(1)
Australia -0.04 80-17(17)
Austria -0.07 98-21(8)
Azerbaijan -2.02 98-19(12)
Bahrain -0.24 10-12(2)
Belgium -0.01 97-21(9)
Bolivia -0.02 00-13(5)
Botswana 0.19 97-97(1)
Brazil -0.03 00-19(8)
British Virgin Islands 12.02 97-97(1)
Bulgaria -0.25 98-16(8)
Burkina Faso 0.04 01-18(8)
Cameroon 1.31 98-19(9)
Canada 0.06 80-20(19)
Cape Verde -1.47 10-16(5)
Chile -0.14 00-21(11)
China 0.00 93-19(14)
Colombia 0.09 97-21(7)
Costa Rica -0.09 14-19(3)
Croatia -0.20 00-21(10)
Curaçao 1.72 03-16(7)
Cyprus -0.13 97-21(11)
Czechia -0.19 95-16(10)
Côte d’Ivoire 1.52 97-00(2)
Denmark 0.15 99-21(10)
Dominican Republic -0.19 07-15(5)
Ecuador -1.17 10-18(6)
Egypt 2.24 99-15(7)
El Salvador 3.32 16-19(4)
Estonia -0.04 00-21(9)
Finland -0.07 98-21(10)
France 0.01 87-21(13)
Germany -0.00 80-21(16)
Greece -0.36 97-21(14)
Guatemala 0.46 07-21(6)
Guinea -0.16 07-10(3)
Honduras 0.27 05-15(5)
Hong Kong 0.09 97-10(6)
Hungary -0.04 97-21(12)
Iceland -0.09 02-13(8)
India 0.10 00-15(7)
Indonesia 0.02 19-19(1)
Iran 1.21 99-16(9)
Ireland -0.03 99-21(11)
Number of periods reported is shown in ().
Country $\theta_{c}$ Periods
Israel 0.58 00-17(7)
Italy 0.14 86-21(11)
Japan -0.10 85-21(13)
Kazakhstan 0.02 97-22(13)
Kuwait 0.00 06-15(5)
Kyrgyzstan 0.02 99-21(10)
Latvia -0.09 00-21(9)
Lithuania -0.10 00-21(10)
Luxembourg -0.00 99-21(12)
Malaysia -0.07 08-15(5)
Malta -0.01 97-21(11)
Mauritius -0.06 17-18(2)
Mexico -0.02 97-21(9)
Moldova -0.20 97-19(12)
Mongolia -0.08 11-21(4)
Morocco 0.49 00-21(8)
Netherlands -0.01 97-21(11)
New Zealand 0.03 97-18(11)
Nicaragua 0.05 10-19(5)
Niger 0.73 98-18(9)
Norway 0.06 82-19(17)
Peru -0.19 10-21(4)
Philippines 1.60 99-22(11)
Poland 0.31 03-21(7)
Portugal -0.05 97-22(10)
Qatar -0.21 04-18(6)
Romania -0.17 98-17(7)
Russia -0.09 99-18(6)
Senegal 1.13 21-21(1)
Serbia 0.25 00-21(11)
Slovakia -0.14 01-21(9)
Slovenia -0.10 99-21(11)
South Africa 0.16 99-21(9)
South Korea 0.07 99-19(8)
Spain 0.07 97-21(12)
Sweden -0.03 81-21(22)
Switzerland -0.03 93-20(14)
Tunisia 3.39 97-06(4)
Turkey 0.03 13-15(2)
USA -0.04 83-20(18)
Ukraine 0.05 97-19(14)
United Kingdom -0.20 81-21(19)
Uzbekistan -0.17 13-21(4)
Vanuatu -0.06 03-06(2)
Venezuela -0.02 00-12(6)
3 Balanced growth
Keynes (1936) taught that households tend to consume a constant proportion of income. Consumption would tend to peak somewhere between a worker’s mid-forties and retirement, then, as income does. Friedman (1957) proposed conversely that households consider expected lifetime income, not current income alone, and meet current consumption needs from accumulated savings as well as current income within that total. Ando and Modigliani (1963) followed this lead by suggesting that younger households tend to save to afford children and retirement, and that older ones dissave to pay consumption costs of both as they arrive.
In any of these scenarios, random distribution of ages could allow phases of more saving in some households to coincide with phases of more consumption in other households. It becomes possible, then, if net saving realizes equal capital growth, for consumption, net saving, capital growth and capital itself all to grow at the same rate. Then net output $Y_{net}$ can grow at that same rate if the capital/output ratio holds constant. This scenario where $C$, $S_{net}$, $K$ and $Y_{net}$ all grow at the same rate is called balanced growth.
We agree that households save more and consume more in phases, that net saving tends to bring equal capital growth at the household scale (see Section 5 below), and that random distribution of phases could explain balanced growth within thrift theory. The observation that the consumption/capital ratio $c^{*}$ does seem to hold more or less constant over scale and time could therefore fit the predictions of thrift theory, at first glance, as well as of free growth theory.
A closer look tells the difference. First, perfectly balanced growth is mathematically possible only where capital growth rate $g(K)$ holds constant. [^5] Because net output/income gives the only source of capital growth, and consumption gives the only avenue of capital depletion after depreciation is recaptured within the value-added chain. That would imply $\Delta g(K)=0$. The record shows no period where $\Delta g(K)$ equals zero or $c^{*}$ equals unity exactly. Even when absolute $\Delta g(K)$ is small, and $c^{*}$ is near unity, Eqs. (4) and (5) still apply in thrift theory. Here too, $\theta_{c}$ is observed to approach zero rather than unity.
Second, for reasons shown in Section 4, we screen out years where absolute $\Delta g(K)$ is less than 0.01. Thus our tests show only results where the predictions of thrift theory and free growth theory differ substantially.
Test results are conclusive. $\theta_{c}$ and the regression of $-\Delta c^{*}$ on $\Delta g(K)$ are both predicted at unity in thrift theory. They are measured respectively at near zero and substantially less than zero. These findings parallel and reinforce our similar ones from data for net saving and consumption reported in Getty and Tkachenko (2024).
4 Screens against small denominators
No measurement is perfect. Small mismeasurements of small denominators, even if numerators were measured accurately, can distort calculations of fractions dramatically. The denominator in calculating $\theta_{c}$ is $\Delta g(K)$, the change in capital growth rate. This change, up or down, is occasionally very small. We control for that risk by screening out years where absolute $\Delta g(K)$ shows as less than $0.01$.
5 Growth by thrift at smaller scales
Free growth theory predicts only at the collective or global scale, meaning the scale of all households and nations taken together. We follow John Rae (1834) in reasoning that some households and communities can grow by saving if others dissave as much. We accept that much growth at the scale of individual households and firms, and possibly even small nations, fits the predictions of thrift theory.
6 Source of data
All our data are drawn from Distributional National Accounts (DINA) shown in the free online database World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. In this paper, we show results for the 90 of those countries which report all information needed for deriving $\theta_{c}$. We find market-value capital $K$ from WID as Capital Wealth (mnweal). We construct consumption as the sum of the consumption expenditures of the government (mcongo) and private expenditures of households and NPISH (mconhn). GDP, which we use only for weighting purposes, is reproduced from GDP (mgdpro).
7 Summary and discussion
Tests of thrift theory from data for consumption support findings from data for net saving in Getty and Tkachenko (2024). This paper refers to that one for logical arguments and policy prescriptions. The finding is that capital growth at very large scales, not small ones, is financed from depreciation investment with no help from net saving. Thus net saving, we conclude, is sacrifice of consumption with nothing in return.
These findings warrant attention. Much of macroeconomics as taught today, whether in Keynesian or anti-Keynesian schools, rests on the assumption that capital growth requires equal net saving. Data show that it requires no net saving. We suggest urgent review of our data and methods, as well as new tests from other data, and study of implications if our findings are confirmed.
[…]